Alice is closing on a house on Aug 13t the buyer owens the property on the day of the closing. the selling price of the home is 350,700. Alice was accepted for a 20 year fixed-rate mortage for 324900 at 5.25% interest. the seller has paid 4478.51. In property taxes for the coming year. How much will Alice owe in prorated taxes and interest?


1
Expert's answer
2021-05-10T15:30:11-0400

Interest portion

Amount of interest for the year"=\$324,900\times5.25\%=\$17,057.25"


Alice portion till August 13"=\$17,057.25\times\frac{225}{365}=\$10,514.74"


Seller portion from August 13"=\$17,057.25\times\frac{140}{365}=\$6,542.51"


Tax portion

Alice portion till August 13"=\$4,478.51\times \frac{225}{365}=\$2,760.73"


Seller portion from August 13"=\$4,478.51\times\frac{140}{365}=\$1,717.78"


Total Alice share"=\$10,514.74+\$2,760.73=\$13,275.47"